OpenStudy (anonymous):

3 --% ×75 = _____ 4 Select one of the options below as your answer:A. .255 B. .5625 C. 2.55 D. 5.625

OpenStudy (freckles):

\[\% =\frac{1}{100}\] \[\frac{3}{4} \cdot \frac{1}{100} \times 75\] \[\frac{3}{4} \cdot .01 \times 75\] \[.75 \cdot .01 \times 75\] So what do you get when you multiply .75 and .01 ?

OpenStudy (freckles):